# EVOLvE: Evaluating and Optimizing LLMs For Exploration

### v1

@@ 6 Related Work
Several prior works have investigated the use of LLMs for decision-making. In one category, numerous studies have deployed LLMs directly as agents in decision-making problems such as games (Yao et al., 2023; Brooks et al., 2024; Shinn et al., 2024; Wang et al., 2023a; Xi et al., 2023). However, fewer works have systematically evaluated LLMs’ capabilities in general decision-making setup, especially in relation to classical concepts in decision-making like exploration. Our work extends the research of Krishnamurthy et al. (2024), who examined LLMs’ exploration capabilities in small-scale MAB tasks. Their findings, which showed positive results only with substantial intervention, are consistent with our broader analysis across both MAB and CB tasks at various scales. Additionally, Mirchandani et al. (2023) also evaluated the ability of LLMs to learn in-context from demonstrations (and demonstrations of improvement) to solve decision-making problems.
The line of research on using LLMs as optimizers faces many similar challenges to in-context decision making, although applied to different tasks. Yang et al. (2024) explored the use of language models as general-purpose optimizers for simple black-box optimization problems, such as prompt optimization, highlighting that a careful balance of exploration and exploitation is critical. Another relevant line of research focuses on in-context learning for decision-making and reinforcement learning (RL) with domain-specific transformers. Laskin et al. (2022) distilled demonstrations from RL algorithms into a transformer and showed that the transformer learns to imitate the RL process to solve new RL tasks. Similarly, Lee et al. (2024) trained transformers with optimal action labels, showing that the model learns to execute posterior sampling for RL (Osband et al., 2013) in-context, which tailors exploration to the underlying distribution of RL tasks. This area has been further studied by Raparthy et al. (2023); Lin et al. (2023). However, these studies focus on domain-specific decision-making, whereas our paper examines general-purpose decision-making capabilities in language models.
@@ A.4 Example of Win-Rate Calculation
A.5 Details on Fitting Regret Function
@@
In the MAB-Hard setting, we can see that more models are having non-zero $\beta$, meaning these models have linear cumulative regret, which indicates lack of in-context self-improvement, because the model is not selecting the optimal arm more and more frequently as $T$ increases. However, we can see that generally Oracle Behavior Fine-Tuned models are doing better.
@@
A.6 Evaluation Implementation Details
@@
A.7 Full List of Models
@@ A.8 Scenario Prompts
A.9 Examples of few-shot demonstrations

### v2

@@ 6 Related Work
Several prior works have investigated the use of LLMs for decision-making. In one category, numerous studies have deployed LLMs directly as agents in decision-making problems such as games (Yao et al., 2023; Brooks et al., 2024; Shinn et al., 2024; Wang et al., 2023a; Xi et al., 2023). However, fewer works have systematically evaluated LLMs’ capabilities in general decision-making setup, especially in relation to classical concepts in decision-making like exploration. Our work extends the research of Krishnamurthy et al. (2024), who examined LLMs’ exploration capabilities in small-scale MAB tasks. Their findings, which showed positive results only with substantial intervention, are consistent with our broader analysis across both MAB and CB tasks at various scales. Mirchandani et al. (2023); Rahn et al. (2024); Felicioni et al. (2024) also evaluated the ability of LLMs to learn in-context and solve bandit-like decision-making problems.
Another relevant line of research focuses on in-context learning for decision-making and reinforcement learning (RL) with domain-specific transformers. Laskin et al. (2022) distilled demonstrations from RL algorithms into a transformer and showed that it learns to imitate the RL process to solve new RL tasks. Similarly, Lee et al. (2024) trained transformers with optimal action labels, showing that the model learns to execute posterior sampling for RL (Osband et al., 2013) in-context. This area has been further studied by Raparthy et al. (2023); Lin et al. (2023); Bai et al. (2023). However, these studies focus on domain-specific decision-making, whereas our paper examines general-purpose decision-making capabilities in language models. Our inference-time algorithm-guided support shares a similar conceptual framework with recent efforts to align LLMs at inference time. These include employing explicit value functions as prefix scorers (Mudgal et al., 2023), and leveraging both implicit and explicit value functions to guide decoding (Liu et al., 2024b). In the realm of algorithm distillation, much of the research on LLMs has concentrated on chain-of-thought (CoT) reasoning (Wang et al., 2023b; Li et al., 2023b), while Gandhi et al. (2024) focused on search and backtracking. Our work focuses on distilling a more complex class of algorithms that involve uncertainty estimation and linear regression.
@@
The motivation of our work is also inspired by the cognitive science literature on human. Prior studies have found that humans balance exploration and exploitation through a mix of directed and random exploration (Wilson et al., 2014), and that these behaviors are supported by distinct neural mechanisms (Daw et al., 2006). Gershman (2018) further decomposes human exploration into algorithmic components. These works provide future directions for evaluating LLM’s behavior in addition to reward Pan et al. (2025). Additionally, efficient exploration also benefits other areas of RL, such as multi-agent collboration (Qu et al., 2024) and preference optimization (Bai et al., 2025).
@@ 7 Conclusion
Software and Data
BanditBench and the inference code have been provided in this GitHub repo and will be updated/monitored regularly: https://github.com/allenanie/EVOLvE. You can install the code with: pip install banditbench.
Acknowledgements
We thank Yaqing Wang for helping some infra on data processing and storage, Aviral Kumar, Yifeng Lu, Haokai Lu, Ice Pasupat, Kefan Dong, Ollie Liu, Deqing Fu for valuable discussions on the project. We thank Ruijie Zheng for feedback on the code base.
@@ A.4 Example of Win-Rate Calculation
A.5 Benchmark Evaluation Cost
We calculate and report the inference time cost for evaluating 30 trials for one agent with the longest contextual representation (which incurs the highest cost). This means, for MAB tasks, we evaluate using Raw History (RH) and for CB tasks, we evaluate using Algorithm-Guided Support (AG). The total cost for MAB Full (over 16 configurations and 32 models) is \$15897.6 for Gemini-1.5 Flash and for CB (2 configurations and 14 models) is \$14491.12 for Gemini-1.5 Flash. We will release all logged experiment data of all models for public analysis and comparison. We also recommend using a subset of our configurations for MAB (HardCore and HardCore+) to reduce evaluation cost. The cost was calculated around 02/2025.
Core HardCore HardCore+ Full MovieBench
Gemini-1.5 Flash \$31.05 \$14.91 \$39.18 \$83.44 \$31.05
gpt-4o-mini-2024-07-18 \$62.10 \$29.83 \$78.36 \$166.88 \$62.10
claude-3-5-haiku-$2$414.33 \$198.97 \$522.64 \$1113.18 \$414.33
Gemini-1.5 Pro \$517.54 \$248.55 \$652.98 \$1390.69 \$517.54
GPT-4o \$1035.07 \$497.11 \$1305.96 \$2781.38 \$1035.07
Claude-3.5-sonnet \$1243.00 \$596.91 \$1567.91 \$3339.53 \$1243.00
Table A3: Maximal inference cost for a single agent/task configuration over 30 trials. Price calculated on 1/30/2025. Core, HardCore and HardCore+ refer to specific subset of the MAB configurations. The exact details are in the code included in the supplementary material. In this paper, we used Full (MAB) and MovieBench (CB) of the benchmark.
A.6 Details on Exploration Optimality Analysis
There are two types of failures one can expect in a bandit problem:
  1. 1.
    Over-exploration on suboptimal choices which results in lower exploration efficiency: over-exploration happens when the algorithm spends too much time exploring suboptimal choices, reducing overall efficiency. This behavior can be quantified using the MinFrac metric (Krishnamurthy et al., 2024), which measures the fraction of pulls allocated to the least-selected arm. An ideal algorithm should exhibit high MinFrac during early exploration (when $T$ is small) and low MinFrac as $T$ increases (indicating effective exploitation).
  2. 2.
    Failure to identify the optimal arm: this occurs when the algorithm struggles to converge on the best option over time. To capture this, we compute the percentage of times an optimal arm is pulled at different time steps (OptFrac). Ideally, this percentage should increase as the process progresses, indicating the model’s ability to self-improve.
We report these metrics at specific time steps over a total of T time steps. For convenience in visualizing the results in a table, we select the 10%, 25%, 50%, 75%, and 100% (final) time steps. There are 1000 steps for the Bernoulli The reported numbers are from Bernoulli Bandit, the metrics are averaged across different configurations.
MinFrac (%) / $t$ Step 100-th 250-th 500-th 750-th 1000-th
UCB 82.3 48.6 27.8 19.6 15.3
Gemma-2B (SH) 0.0 1.0 0.5 0.4 0.3
Gemma-2B (AG) 0.0 0.6 1.1 1.9 2.7
Gemma-9B (SH) 6.9 11.2 16.9 15.4 16.5
Gemma-9B (AG) 5.8 11.8 17.3 19.6 22.9
Gemini-1.5 Flash (SH) 10.2 4.2 2.1 1.4 1.1
Gemini-1.5 Flash (AG) 11.3 4.5 2.3 1.5 1.1
Gemini-1.5 Pro (SH) 79.0 40.0 20.6 13.9 10.5
Gemini-1.5 Pro (AG) 73.8 37.1 18.9 12.7 9.5
Table A4: Over-exploration Rate (MinFrac): MinFrac measures the fraction of pulls allocated to the least-selected arm. We show the measures on the $t$-th step along the progression of exploration.
OptFrac (%) / $t$ Step 100-th 250-th 500-th 750-th 1000-th
UCB 32.7 49.4 58.7 62.6 65.0
Gemma-2B (SH) 10.1 10.3 10.2 10.1 10.1
Gemma-2B (AG) 12.8 12.3 12.4 12.2 12.2
Gemma-9B (SH) 5.8 6.7 7.4 7.8 8.0
Gemma-9B (AG) 6.6 7.3 6.7 6.6 6.6
Gemini-1.5 Flash (SH) 9.3 10.1 10.4 10.6 10.7
Gemini-1.5 Flash (AG) 14.4 15.6 16.3 16.6 16.8
Gemini-1.5 Pro (SH) 15.1 19.1 21.8 22.9 23.5
Gemini-1.5 Pro (AG) 11.9 17.7 21.6 23.1 23.9
Table A5: Fraction of Pulls on the Optimal Arm (OptFrac): OptFrac measures the fraction of pulls overall on the optimal arm over. We show the measures on the $t$-th step along the progression of exploration.
A model could, in theory, achieve high performance by consistently choosing the optimal arm—even if it rarely explores—by chance. To address this, we include an analysis of the model’s exploration behavior using a metric called OptFrac, which measures how often the optimal arms are selected. As shown in Table A5, UCB steadily increases its OptFrac over time (32.7% $\rightarrow$ 49.4% $\rightarrow$ 58.7% $\rightarrow$ 62.6% $\rightarrow$ 65.0% over 1000 steps), indicating a growing focus on the optimal arm. In contrast, Gemini-1.5 Flash remains largely flat (9.3% $\rightarrow$ 10.1% $\rightarrow$ 10.4% $\rightarrow$ 10.6% $\rightarrow$ 10.7%), suggesting that it does not significantly shift its behavior toward the optimal arm. This supports our claim that the model does not accidentally achieve optimal performance by randomly selecting the best arm without meaningful exploration.
We then analyze exploration dynamics in more detail, such as whether the model engages in random or directed exploration. In our analysis, we include a metric called MinFrac, which measures the fraction of pulls allocated to the least-selected arm. This captures the extent to which the model explores less-visited options, which can be understood as a form of “directed exploration.” Ideally, this value should be high early on (indicating strong directed exploration), and then decrease as the model gains experience and focuses on better-performing arms. As shown in Table A4, UCB exhibits this expected trend, with MinFrac values decreasing over time: 82.3% $\rightarrow$ 48.6% $\rightarrow$ 27.8% $\rightarrow$ 19.6% $\rightarrow$ 15.3%. In contrast, Gemini-1.5 Flash starts with a much lower MinFrac and declines rapidly (11.3% $\rightarrow$ 4.5% $\rightarrow$ 2.3% $\rightarrow$ 1.5% $\rightarrow$ 1.1%), suggesting it lacks meaningful directed exploration from the outset.
A.7 Details on Fitting Regret Function
@@
In the MAB-Hard setting, we can see that more models are having non-zero $\beta$, meaning these models have linear cumulative regret, which indicates lack of in-context self-improvement, because the model is not selecting the optimal arm more and more frequently as $T$ increases. However, we can see that generally Optimal Behavior Fine-Tuned models are doing better.
@@
A.8 Evaluation Implementation Details
@@
A.9 Full List of Models
@@ A.10 Scenario Prompts
A.11 Examples of few-shot demonstrations